40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 26 20:12:18 2010, kiwamu, Update, SUS, OSEMs on PRM DSC_2279.JPG
    Reply  Tue Jul 27 21:38:21 2010, kiwamu, Update, SUS, installation of in-vac optics DSC_2284.JPGDSC_2291.JPG
       Reply  Wed Jul 28 17:32:37 2010, kiwamu, Update, SUS, installation of in-vac optics 
          Reply  Thu Jul 29 10:39:43 2010, kiwamu, Update, SUS, Re: installation of in-vac optics 
Message ID: 3294     Entry time: Mon Jul 26 20:12:18 2010     Reply to this: 3302
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: OSEMs on PRM 

 [Alberto and Kiwamu]

We installed the OSEMs to the new PRM.

As I wrote down on the elog (see here)  today's mission was to install the OSEMs to the PRM.

After putting them on the tower we adjusted the readout offsets by sliding the OSEMs so that they can stay in the linear sensing ranges. 

Now all of the OSEMs have almost good separation distances from the PRM.

In the attached picture you can see the OSEMs installed on the PRM tower ( middle: PRM tower, left: BS tower)


(what we did)

 1. moved the PRM tower close to the door so that we could easily access the PRM.

 2. leveled the table by putting some weights and confirmed the level by a  bubble level tool.

     - We must level the table every time when we set / adjust any OSEMs,  otherwise the readout voltages of  the OSEMs vary every time due to the tilted table.

 3. released the PRM by loosing the earthquake stops

 4. put the OSEMs with approximately right separation distances from the PRM.

      -  At this phase we can see the readout of the OSEMs, which were oscillating freely because we still didn't enable the damping.

        -  The OSEM positions were checked by looking at useful notes on the wiki (see here).

 5. turned on the damping servo of the OSEMs

       - Without changing any gains, it worked well. 

      - Then we could see stable readouts of the OSEMs which didn't show any oscillations in turn because of the damping.

 6. checked the level of the table again

 7. set each of the OSEM readouts to the half of its maximum value by sliding their positions slightly.

      - The readout offsets were at almost the half value within +/- 100 mV accuracy (this was the best accuracy we could adjust by our hands)

 8. screwed down the earthquake stops to lock the PRM.

      - Now the damping is off.

 9. closed the door

 


(to be done)

 *  Putting the PRM tower back to the designed place

 *  Installation of the pick off mirror

 *  Arrangement of the PZT mirror

Attachment 1: DSC_2279.JPG  2.450 MB  | Hide | Hide all
DSC_2279.JPG
ELOG V3.1.3-